OpenStudy (anonymous):

Find the standard form of the equation of the ellipse with the given characteristics. vertices: (- 8, - 5), (- 8, 9) minor axis of length: 4

OpenStudy (dumbcow):

what do you have so far?

OpenStudy (dumbcow):

there are 3 things you need: center of ellipse a^2 b^2

OpenStudy (anonymous):

center is (-8, 2) is the center?

OpenStudy (dumbcow):

yep

OpenStudy (dumbcow):

\[\frac{(y-2)^{2}}{a^{2}}+\frac{(x+8)^{2}}{b^{2}}=1\]

OpenStudy (jdoe0001):

"b" component is given pretty much minor axis is 4, "b" is half that b = 2 from (- 8, - 5) to (- 8, 9) moving up/down, it goes from -5 up to 9, 14 units for the major axis, half of that is "a" a = 7